what is a line perpendicular to y=-3/12x+2 and passes through the x-axis at 3?
what is a line perpendicular to y=-3/12x+2 and passes through the x-axis at 3?
Y=-3/12 x +2 when x=3
=-3/12*3 +2
=-9/12 +2
=-9/12 + 24/12
=24/12-9/12
=15/12
=5/4
Y = 5/4

Answer by chiexpert(48)

what is a line perpendicular to y=-3/12x+2 and passes through the x-axis at 3?
Y=-3/12 x +2 when x=3
=-3/12*3 +2
=-9/12 +2
=-9/12 + 24/12
=24/12-9/12
=15/12 divisible by 3
=5/4
Y = 5/4